OpenStudy (anonymous):

Select all the true statements from below: A. 6, 8 and 10 are the lengths of the 3 sides of a right triangle. B. 10, 12 and 15 are the lengths of the 3 sides of a right triangle. C. The square root of 225 is larger than 17. D. 10, 24 and 26 are the lengths of the 3 sides of a right triangle. E. 9, 15 and 17 are the lengths of the 3 sides of a right triangle.

OpenStudy (anonymous):

@sammixboo

OpenStudy (gebooors):

Try to calculate, is Pythagoras theorem A^2 + b^2 = c^2 valid

OpenStudy (anonymous):

so like 6^2+8^2+10^2

OpenStudy (anonymous):

oh...so 6^2 + 8^2 = 10^2

OpenStudy (anonymous):

Thatd be 6x6 + 8x8 = 10x10

OpenStudy (anonymous):

A is wrong B is wrong D is wrong E is wrong so the only right one is C? @Gebooors

OpenStudy (gebooors):

Recalculate A 6^2 = 36, 8^2 = 64 and sum is 100. What is 10^2?

OpenStudy (anonymous):

100

OpenStudy (gebooors):

And square root of 225 is 15 in C

OpenStudy (anonymous):

So A is the only right answer?